2) Could you calculate the component cost of equity for a stock with nonconstant expected growth rates in dividends if you didn't have the information necessary to compute the component cost using the CAPM? Why or why not??
Answer #1

Yes you could calculate the cost of equity from non constant expected growth rates. This is because cost of equity can be calculated for non constant growth using dividend discount model for non constant growth.
D1 =D*(1+g1),
D2=D1*(1+g1)*(1+g2)..................Dn=Dn-1*(1+gn). Where d1, d2 .....dn are dividends in year 1, 2 and....n years respectively.
If Price of stock is given . Then required rate or financial calculator can be used to calculate cost of equity. In this cost of equity includes capital gain and dividend yield.